Suppose that a companies production function is given by: f(L;K) = (10K^3L^2)/(L+K)

a) Does this production function exhibit increasing, constant, or decreasing returns to scale? Algebraically justify your answer.

b) If there is a wage of 10 and a rental rate of capital of 1, then find the company's expansion path.
Answer #1

e) FLLIK) = 10K² L² Ltk FCAL, dk) = lorax) 3 (AL) alt dk) 3 lok 3 d2 (2 A[17) 15 Lok??) Ltk = t ( 10k??) | LK d4 ECL, K) = ex6) - E(lik) = 10x3 22 ( 2) MP2 = OF = (L+K) (20k3L)- 10x3,² (170) 22 . Catk) 2 0 - Jok 31 [ 2(2+X) - 2) 1L +x)2 = 10k3L ( 22+Lok² L (L+2K) & stor (tkot tok ²2² (32Z2x) K.KxLL +2K) xrx.2 (32+2K) KL[+2x) 2 (32 +21) KL+zk² 322 2KL Expansion MRTS path is